OpenStudy (anonymous):

2 (x+4) = 12-4x I can't seem to solve it. Whenever I try to solve it I get stuck and get 2=x as my answer but when I plug it in the equation it doesn't equal both sides. Help would be appreciated thanks c:

OpenStudy (anonymous):

2 (x+4) = 12-4x 2x + 8 = 12 -4x +4x +4x 6x + 8 = 12 -8 -8 6x = 4 /6 /6 x = 4/6=2/3

OpenStudy (anonymous):

2x+8=12-4x 2x+4x=12-8 6x=4 x=4/6

OpenStudy (mathstudent55):

Distribute on left side: 2x + 8 = 12 - 4x Add 4x to both sides; Subtract 8 from both sides: 6x = 4 Continue...

OpenStudy (anonymous):

OH. I see what I did now to get that answer. Thank you all c: